John is athird of his father's age.In ten years time,he will be ahalf of his father's age then.How old is john
Answer by ikleyn(52781)

.
John is athird of his father's age. In ten years time, he will be ahalf of his father's age then.How old is john
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

F = 3J            (1)    (J = John's present age; F = father's present age)
F+10 = 2(J+10).   (2)    ( "In ten years time, John will be ahalf of his father's age then" )

Substitute (1) into (2). You will get

3J + 10 = 2(J + 10),   or

3J + 10 = 2J + 20  --->  J = 20 - 10  --->  J = 10.

Answer.  John is ten years now. His father is 3*10 = 30 years now.

For age problems see the lessons
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.